# Signage-Aware Exploration in Open World using Venue Maps

### v1

@@ Abstract
Current exploration methods struggle to search for shops in unknown open-world environments due to a lack of prior knowledge and text recognition capabilities. Venue maps offer valuable information that can aid exploration planning by correlating scene signage with map data. However, the arbitrary shapes and styles of the text on signage, along with multi-view inconsistencies, pose significant challenges for accurate recognition by robots. Additionally, the discrepancies between real-world environments and venue maps hinder the incorporation of text information into planners. This paper introduces a novel signage-aware exploration system to address these challenges, enabling the robot to utilize venue maps effectively. We propose a signage understanding method that accurately detects and recognizes the text on signage using a diffusion-based text instance retrieval method combined with a 2D-to-3D semantic fusion strategy. Furthermore, we design a venue map-guided exploration-exploitation planner that balances exploration in unknown regions using a directional heuristic derived from venue maps with exploitation to get close and adjust orientation for better recognition. Experiments in large-scale shopping malls demonstrate our method’s superior signage recognition accuracy and coverage efficiency, outperforming state-of-the-art scene text spotting methods and traditional exploration methods.
@@
Scene Understanding, Scene Text Spotting, Exploration-Exploitation, Path Planning.
@@ I Introduction
Searching for shops using their signage information in unknown open-world environments poses a significant challenge for robots due to a lack of prior knowledge. In contrast, humans can effortlessly navigate to the goal landmarks by leveraging the venue maps in daily life. Imagine walking through a shopping mall: we locate ourselves on the venue maps based on actual observations and then plan paths on the maps to reach our destinations. Significant efforts have been made to correlate the visual observations with venue maps for global localization [1, 2, 3, 4, 5], or to discover non-metric heuristic derived from maps for kilometer-scale navigation [6, 7]. However, these approaches primarily focus on using geometric and semantic information from the venue maps, overlooking valuable cues from landmark names on the signage. This oversight is crucial for accurately identifying locations and reducing ambiguity in localization [8]. Moreover, they do not consider online scene understanding and exploration. Current exploration methods generally incorporate geometric information [9, 10] or object-level scene semantics [11, 12, 13, 14, 15, 16], but they often neglect text-level semantics, limiting their effective use of venue maps. While some efforts detect room names as cues in office environments [17, 18, 19], they do not handle signage recognition. We argue that a signage-aware exploration method is essential to bridge the gap between searching for shops using venue maps and exploration methods.
@@
Figure 1: We propose to leverage the venue map as the prior knowledge to facilitate shop searching in unknown open-world environments. In the figure, given a venue map, the robot recognizes the text on the signage of the nearby restaurant and localizes itself on the map. Then the robot knows the direction towards next landmark, which is called ‘Briketenia,’ and searches for it.
@@
However, implementing such a signage-aware exploration method poses two main challenges. First, recognizing shop names on signage in open-world environments is difficult as they are presented in arbitrary shapes and styles, along with multi-view semantics inconsistencies. Existing works in text recognition for visual localization [1, 2] or place recognition [20, 21, 22, 23, 24] often rely on the OCR models trained on closed-set datasets, which struggle to recognize signage effectively in open-world scenarios. To overcome this, an online signage understanding, i.e., text-centric scene understanding method, and an active signage perception policy are essential. Second, using venue maps to guide exploration planning is challenging due to their inconsistent scales and distortions compared to real-world situations. We propose leveraging the directional relationship between landmarks on venue maps as a heuristic to guide exploration without relying on metric information.
@@
This paper presents a signage-aware exploration system that leverages venue maps to facilitate searching for shops in the unknown open world. Our method performs a human-like retrieval process rather than directly employing a conventional closed-set text recognition approach. It identifies the landmark name on the signage by finding the closest match within a pre-established text set extracted from venue maps based on visual similarity. This allows the closed-set detectors to adapt to open-world situations without fine-tuning. We also project multi-view 2D text regions into 3D space and fuse the corresponding features to enhance recognition accuracy. Furthermore, we design a venue map-guided exploration-exploitation planner that enables the robot to search for the signage using a directional heuristic. Once a signage candidate is detected from a distance, the system switches to exploit known spaces to approach the signage and adjust the robot’s view to recognize the signage clearly. By balancing exploration and exploitation, we achieve both high signage recognition accuracy and coverage efficiency. Our main contributions to this work are summarized as follows:
@@
  • A venue map-guided exploration-exploitation planner for efficiently covering the signage in the open world and meanwhile establishing a signage map.
  • Evaluation experiments in real-world shopping malls have substantiated our method’s efficacy and robustness.
Figure 2: Overall framework of the proposed method. Our method first pre-builds a topological graph based on the spatial relationship between landmark names on a given venue map (Sec. IV-A). Then, given the RGB-D image and point cloud at each timestep, our method first employs a scene text detector to detect the text instances on the observed signage. Next, we project the 2D text regions to 3D space and perform 3D fusion if they can be merged with the previously detected text instances. Then, we introduce a diffusion-based text instance retrieval method to correlate the detected text with the text set on the venue map (Sec. IV-B). Once localized on the venue map, the next landmark goal is inferred by the topological planning and used to guide the selection of frontiers. The detected frontiers and the viewpoints generated at the signage are jointly fed to the planner for a decision on the next best pose. Throughout the process, our approach balances exploration and exploitation to improve signage recognition accuracy and coverage efficiency (Sec. IV-C).
@@ II-A Discovering the Potential of Venue Maps
Venue maps, satellite maps or OpenStreetMaps (OSM), are commonly encountered in our daily lives, offering valuable and readily accessible information for human and robot navigation. Existing works primarily study estimating the global pose $(x,y,\theta)$ on the 2D map’s coordinate using visual observation over the latent space [1, 2, 3, 4, 5], such as SNAP [5], which aligns the query image with a fuse of ground-level and overhead imagery over a learned 2D neural space, or inducing the potential heuristics from the map to achieve long-range point-goal navigation in open-world scenarios [6, 7]. ViKiNG [6] learns a heuristic function to predict whether future steps lie in the paths towards the goal, and S2MAT [7] utilizes a user-specified GPS sequence to guide the navigation. As for leveraging the landmark names on the venue maps, [1] introduces the shop names into the global localization and formulate the problem as inference in a Markov random field, and [2] employs a particle filter into text detection for pose estimation. However, correlating shop names with the venue maps to facilitate exploration planning has not been studied.
@@ II-B Semantic-aware Exploration
Semantic-aware exploration methods that focus on the semantics in the scene have arisen in recent years [11, 12, 13, 14]. They integrate scene understanding methods with traditional planning, seeking both mapping efficiency and semantic coverage. SARHP [14] proposes an object-centric planner for optimizing the semantic gain of the exploration paths. 3D scene graph methods that focus on the objects and their topological relations to enable high-level reasoning for robotic navigation and manipulation have also been developed, such as ConceptGraphs [15] and HOV-SG [16]. They segment the object masks via a proposal network such as SAM [25] and extract the language-aligned features of the cropped regions via CLIP [26] to support open-vocabulary query, which assumes a sequence of RGBD trajectory is pre-collected manually or by any exploration method and focus on establishing scene representation on it. Different from the above works that concentrate on object-level semantics, and also distinct from studies targeting simple text cues, e.g., the room names in office environments [17, 18, 19], our work focuses on incorporating the text information on signage into exploration planning.
@@ II-C Text-based Place Recognition and Scene Text Spotting
Text features have garnered significant attention in visual place recognition (VPR) task [20, 21, 22, 23], which employs text-based descriptors to facilitate matching between a query image and a database of scene images, due to their robustness to the appearance changes and distinctive features in different places. Recently, TextSLAM [24] develops a novel SLAM system tightly coupled with text features to build a 3D volumetric map and enhance the pose estimation accuracy. On the other hand, scene text spotting (STS) task that detects and recognizes the text in natural scenes have drawn great attention recently, particularly spotting the text on the signage, which is more difficult in terms of styles and fonts. ESTextSpotter [27] addresses this by explicitly synergying the text detection and recognition modules, achieving state-of-the-art performance.
Despite using the optical character recognition (OCR) models such as PP-OCRv3 [28] and STS models like ESTextSpotter [27], the signage recognition results are often too noisy due to the partial observation and multi-view inconsistency issues. [1] adopts bag-of-K gram to mitigate the former issue, which requires predicting a 10000-dimensional classification vector and is still unsuitable for square characters. To make a general solution, we conduct similarity matching between the observed signage and the text set of venue maps to improve the signage recognition accuracy without fine-tuning. This is similar to the scene text retrieval (STR) task but with different motivations. Prior works in STR convert the queried text to an image via font rendering and perform image-level similarity matching to localize and retrieve all the text instances from a pre-established image gallery [29, 30, 31]. In contrast, we aim to improve the signage recognition accuracy and use a text-diffusion model for text-to-image conversion.
@@ III Problem Formulation
We aim to design an online signage-aware exploration method that leverages the venue map for searching for shops in unknown large-scale human-oriented environments such as shopping malls. We model the environment as a 3D space $V\in\mathbb{R}^{3}$, where the points $\mathbf{v}\in V$ are classified into occupied, free, and unknown with the associated probability $p_{o}(\mathbf{v})$. The environment also contains $N$ static landmarks that can be shops or restaurants identified by their names $\mathcal{T}$. The venue map M (see Fig. 1) portrays all the landmarks and the inaccurate topological relationship $\mathcal{G}$ among them. Note that we do not necessitate any geometric information or semantic layouts since some maps may not contain them or require additional segmentation that limit the generalization of the method. The planner takes as inputs the venue map M, RGB image $o$, depth $d$, and LiDAR points $\mathcal{X}$ at each timestep, outputs next best pose $c\in\mathcal{F}\cup\mathcal{V}$ among the sampled frontiers $f\in\mathcal{F}$ and viewpoints $v\in\mathcal{V}$, while concentrating on the signage in the scene. During the process, a signage map $\mathcal{M}$ is established online, where the recognized signage text $\tau\in\mathcal{T}$ and the corresponding located regions are stored.
@@ IV Methodology
Our method consists of three parts: 1) topological planning on venue maps, 2) signage understanding, and 3) exploration-exploitation planning. The framework is illustrated in Fig. 2.
@@ IV-A Topological Planning on Venue Maps
We first pre-build a topological graph $\mathcal{G}$ on the venue maps whose nodes are the landmark names detected by an OCR model and edges connect two nodes if their Euclidean distance on the map is lower than a threshold $\gamma_{map}$. For the isolated subgraphs, we iteratively merge them at the nearest node pairs. As such, we solve a travel salesman problem (TSP) to obtain a landmark route $g_{1:K}$. During online exploration, our method searches and recognizes the signage in the scene step-by-step to handle long-horizon planning [32].
Figure 3: The pipeline of signage understanding. (a) At the offline stage, we use a text-diffusion model to render the landmark names on the generative images. (b) At the online stage, we project the detected text images to 3D space and fuse the features with previous observations of the same signage. Then we retrieve the most similar offline images compared to the detected images to determine the labels, which are projected onto a signage map. All the text features are extracted by a scene text spotter $\phi$.
@@ IV-B Signage Understanding
We online detect and recognize the signage using the proposed signage understanding method. As such, the robot is globally localized on the venue maps. Meanwhile, we construct a signage map $\mathcal{M}$ for downstream querying and planning, which stores the recognized signage text $\tau$ and the corresponding located regions. In practice, we assign the clustered contours of the online map that is nearest to the signage as its corresponding landmark’s located regions.
@@
Diffusion-based Text Instance Retrieval: Current OCR or STS models are limited by their closed-set training that cannot handle signage recognition with arbitrary shapes, styles, and multi-view inconsistency issues in the open world (see Fig. 6). To address this, we leverage the text set of the venue maps to perform visual similarity matching that enables the closed-set detectors to adapt to the open world without fine-tuning. Specifically, we employ a two-stage process for signage recognition. At the offline stage, we utilize an off-the-shelf multilingual text-diffusion model, AnyText [33], to convert all the landmark names $\mathcal{L}$ to the text-rendered signage images set $\{o^{\prime}\}$. Then, we adopt ESTextSpotter [27], denoted as $\phi$, to detect text regions and only extract the text features before the recognition heads from these offline-generated images without using the recognition results, thereby creating a prior features pool $\mathcal{D}_{f}$. During online exploration, we utilize the same spotter $\phi$ to detect the signage from the current image observation $o_{t}$ and extract its features $f_{t}$. We then calculate the cosine similarity $s_{\phi}(o_{t},o^{\prime}_{j})$ between the detected text features and those of the generative text images to find the most similar pair. In practice, we also enhance the retrieval process by narrowing down the matching scope to a subset of $h$-hop neighbouring to the previous recognized landmark $\tau_{t-1}$ on the venue maps based on the robot’s real-time location, resulting in better accuracy and faster speed:
$s_{\phi}(o_{t},o^{\prime}_{j})=cos(\phi(o_{t}),\phi(o^{\prime}_{j})),\ j\in Neighbour(\mathcal{G},\tau_{t-1},h).$ (1)
@@
As such, the detected scene text $\tau_{t}$ is recognized and matched to the most similar landmark name at the image level, i.e., $\tau_{t}=\mathcal{T}_{arg\max s_{\phi}(o_{t},o^{\prime}_{j})}$, with a recognition score $s_{\phi}(o_{t},o^{\prime}_{j})$.
@@
2D-to-3D Fusion: To mitigate the noisy text recognition from single-view 2D images, we propose adopting a 2D-to-3D instance fusion strategy [34] to enhance recognition robustness using multiple observations. During exploration, the signage map serves as a 3D text bank $\mathcal{B}$ that maintains the merged 3D text instances $R^{3D}_{j}$. At each frame $o_{t}$, we first project the detected 2D text regions $\{r_{i}^{2D},..,r_{k}^{2D}\}$ to 3D regions $\{r_{i}^{3D},..,r_{k}^{3D}\}$ using depth $d_{t}$, pose $P_{t}$ and camera intrinsics $C$. Subsequently, for each new 3D text region $r_{i}^{3D}$ and each existing 3D text instance $R^{3D}_{j}$ in the bank, if their cosine similarity $s(f^{2D}_{i},f^{3D}_{j})$ exceeds a predefined threshold $\gamma_{s}$ and the 3D distance between their centroid points $d(r_{i}^{3D},R_{j}^{3D})$ is lower than a threshold $\gamma_{d}$, then we fuse two instances by aggregating centroid points and averaging over the features:
$R^{3D}_{j}:=R^{3D}_{j}\cup r_{i}^{3D}~{}\text{and}~{}f_{j}^{3D}:=(nf_{j}^{3D}+f_{i}^{2D})/(n+1).$ (2)

### v2

@@ Abstract
Current exploration methods struggle to search for shops or restaurants in unknown open-world environments due to the lack of prior knowledge. Humans can leverage venue maps that offer valuable scene priors to aid exploration planning by correlating the signage in the scene with landmark names on the map. However, arbitrary shapes and styles of the texts on signage, along with multi-view inconsistencies, pose significant challenges for robots to recognize them accurately. Additionally, discrepancies between real-world environments and venue maps hinder the integration of text-level information into the planners. This paper introduces a novel signage-aware exploration system to address these challenges, enabling the robots to utilize venue maps effectively. We propose a signage understanding method that accurately detects and recognizes the texts on signage using a diffusion-based text instance retrieval method combined with a 2D-to-3D semantic fusion strategy. Furthermore, we design a venue map-guided exploration-exploitation planner that balances exploration in unknown regions using directional heuristics derived from venue maps and exploitation to get close and adjust orientation for better recognition. Experiments in large-scale shopping malls demonstrate our method’s superior signage recognition performance and search efficiency, surpassing state-of-the-art text spotting methods and traditional exploration approaches. Project website: sites.google.com/view/signage-aware-exploration.
@@
Autonomous Agents, Semantic Scene Understanding, Mapping, Planning under Uncertainty
@@ I Introduction
Humans can efficiently navigate and explore a mall to search for shops or restaurants using a 2D venue map provided by the mall or Google Maps, even if the venue map is non-metric, illustrating only relative relationships between the different landmarks. In contrast, robots struggle to search for landmarks (here we refer to shops or restaurants) in unknown environments due to a lack of scene priors. However, this capability is fundamental for various applications, such as delivery, tour guidance, and inspection. Recently, significant efforts have been made to improve robots by mimicking human behaviors, e.g., correlating the visual observations with venue maps for global localization [1, 2, 3, 4, 5], or by discovering non-metric heuristic derived from maps for kilometer-scale navigation [6, 7]. However, these approaches primarily focus on using geometric and semantic information from the venue maps but overlook the landmark names portrayed on the maps and the corresponding signage displaying the names in the scenes. Moreover, they often neglect online scene understanding and exploration. On the other hand, current exploration methods generally integrate geometric information [8, 9] or object-level scene semantics [10, 11, 12, 13, 14, 15], but rarely incorporate text-level semantics, limiting their effective use of venue maps. While some efforts detect direction signs and arrows [16, 17, 18], and room names [19, 20] as cues in public scenarios, they do not address the recognition of landmark signage. Observing that humans exploit these landmark names on the venue maps by matching them to the signage in the real-world scene, we argue that a signage-aware exploration method is essential to improve a robot’s ability to search for landmarks using 2D (non-metric) venue maps.
@@
Figure 1: We propose to leverage the textual information in a venue map to facilitate shop searching in unknown open-world environments. The robot localizes itself in the environment by recognizing and matching the texts on a sign to the venue map. Then the robot plans a direction to the next landmark ‘Briketenia’.
@@
However, implementing such a system poses the following key challenges. First, recognizing the landmark names on the signage in an open-world environment is inherently difficult. Existing works in text recognition for visual localization [1, 2] or place recognition [21, 22, 23, 24, 25] often rely on optical character recognition (OCR) models trained on closed-set shapes and styles, which struggle to recognize signage effectively in open-world scenarios, where signage can exhibit diverse shapes, styles, and multi-view inconsistencies. Second, the accuracy of signage recognition also depends significantly on the robot’s distance and orientation relative to the signage during movement, which requires an active perception policy to obtain clear observations. Third, using venue maps to guide exploration planning is challenging due to their inconsistent scales and distortions compared to real-world situations.
@@
This paper presents a robotic exploration approach that leverages the signage in an unknown environment and the corresponding (non-metric) venue map to facilitate exploration and searching for landmarks. To address the first challenge, instead of relying on an OCR model trained to recognize texts in limited shapes and styles, we propose a human-like retrieval process based on the appearance of signage. It detects the texts on the signage and matches them with a set of pre-generated signage images based on the landmark names extracted from the given venue map. This allows a closed-set detector to adapt to open-world situations without fine-tuning. We also project multi-view 2D text regions into 3D space and fuse the corresponding features to enhance recognition performance. Furthermore, we design a venue map-guided exploration-exploitation planner that enables the robot to search for the signage corresponding to the landmarks using a directional heuristic. Once a sign candidate is detected, the system will exploit the known space to approach the sign and adjust the robot’s view to faithfully recognize the sign. By balancing exploration and exploitation, we achieve both high signage coverage rates and search efficiency. Our main contributions to this work are summarized as follows:
@@
  • A venue map-guided exploration-exploitation planner for efficiently covering the signage in the open world while grounding the signs in the map.
  • Evaluation experiments in real-world scenes have substantiated our method’s efficacy and robustness.
@@ II-A Discovering the Potential of Venue Maps
Venue maps, satellite maps, or OpenStreetMaps (OSM), are commonly encountered in our daily lives, offering valuable and readily accessible information for human and robot navigation. Existing works primarily study estimating the global pose $(x,y,\theta)$ on the 2D map’s coordinate using visual observation over the latent space [1, 2, 3, 4, 5], such as SNAP [5], which aligns the query image with a fuse of ground-level and overhead imagery over a learned 2D neural space, or inducing the potential heuristics from the map to achieve long-range point-goal navigation in open-world scenarios [6, 7]. ViKiNG [6] learns a heuristic function to predict whether future steps lie in the paths towards the goal, and S2MAT [7] utilizes a user-specified GPS sequence to guide the navigation. As for leveraging the landmark names on the venue maps, [1] introduces the shop names into the global localization and formulates the problem as inference in a Markov random field, and [2] employs a particle filter into text detection for pose estimation. However, correlating shop names with the venue maps to facilitate exploration planning has not been studied.
Figure 2: Overall framework. Our method first constructs a topological graph on a given venue map (Sec. IV-A). Then, given the RGB-D image, the proposed signage understanding method recognizes the texts on the signage and correlates them with the text set of the venue map (Sec. IV-B). Once localized on the venue map, the next landmark goal is inferred to guide the selection of frontiers. Our system balances exploration and exploitation to improve both signage coverage rates and search efficiency during the process (Sec. IV-C).
@@ II-B Semantic-aware Exploration
Semantic-aware exploration methods that focus on the semantics in the scene have arisen in recent years [10, 11, 12, 13]. They integrate scene understanding methods with traditional planning, seeking both mapping efficiency and semantic coverage. SARHP [13] proposes an object-centric planner for optimizing the semantic gain of the exploration paths. 3D scene graph methods that focus on the objects and their topological relations to enable high-level reasoning for robotic navigation and manipulation have also been developed, such as ConceptGraphs [14] and HOV-SG [15]. They segment the object masks via a proposal network, such as SAM [26], and extract the language-aligned features of the cropped regions via CLIP [27] to support open-vocabulary query, which assumes a sequence of RGB-D trajectory is pre-collected manually or by any exploration method. Different from the above works that concentrate on object-level semantics, some studies target utilizing the direction signs and arrows [16, 17, 18] or room names [19, 20] as navigation cues. Maye et al. [16] introduce Hierarchical Implicit Shape Models for robustly recognizing different kinds of signs with only unsupervised training. Our work focuses on incorporating landmark names on the signage into exploration planning.
@@ II-C Text-based Place Recognition and Scene Text Spotting
Text features have garnered significant attention [21, 22, 23, 24] in visual place recognition (VPR) tasks, which employs text-based descriptors to facilitate matching between a query image and a database of scene images, due to their robustness to the appearance changes and distinctive features in different places. TextSLAM [25] develops a novel SLAM system tightly coupled with text features to build a 3D volumetric map and enhance pose estimation accuracy. On the other hand, scene text spotting (STS) task that detects and recognizes the texts in natural scenes have also drawn great attention. ESTextSpotter [28] achieves state-of-the-art performance by explicitly synergizing the text detection and recognition modules.
Despite using the OCR models such as PP-OCRv3 [29] and STS models [28], the signage recognition results are often too noisy due to the partial observation and multi-view inconsistency issues. Wang et al. [1] adopt a bag-of-N-gram model to mitigate the former issue, which requires training a 10000-dimensional classifier to predict whether each composition of the alphabet with a maximum length of N is the substring of the recognized texts and is still infeasible for square-shaped character languages that contain strokes. To make a general solution, we conduct similarity matching between the observed signage and the text set of venue maps to improve the signage recognition performance without fine-tuning. This is similar to the scene text retrieval (STR) task but with different motivations. Prior works in STR convert the queried text to an image via font rendering and perform image-level similarity matching to localize and retrieve all the text instances from an image gallery [30, 31, 32]. In contrast, we improve the signage recognition performance and use a text-diffusion model for text-to-image conversion.
@@ III Problem Formulation
[…]
@@ IV-A Topological Planning on Venue Maps
We first pre-build a topological graph $\mathcal{G}$ on the given venue map whose nodes are the landmark names detected by an OCR model and whose edges connect two nodes if their Euclidean distance on the venue map is lower than a threshold $\gamma_{map}$. For any pair of isolated subgraphs, we merge them with their nearest nodes. As such, we solve a travel salesman problem (TSP) to obtain a landmark route $g_{1:K}$. During online exploration, our method searches for the remaining landmarks sequentially to handle long-horizon planning [33].
Figure 3: The pipeline of signage understanding. (a) In the offline stage, we use a text-diffusion model to render the landmark names on the generative images. (b) In the online stage, we project the detected text images to 3D space and fuse the features with those of multi-view images. Then we retrieve the most similar offline images compared to the detected images as the results, which are projected onto a signage map. All text features are extracted by a scene text spotter $\phi$.
@@ IV-B Signage Understanding
We online detect and recognize the signage using the proposed signage understanding method, which enables the robot to globally localize itself on the venue maps. Meanwhile, we construct a signage map $\mathcal{M}$ for downstream querying and planning, which stores the recognized landmark names $\tau$ on the signage and the corresponding located regions. In practice, we assign the clustered contours of the online map that are nearest to the signage as its corresponding region.
@@
Diffusion-based Text Instance Retrieval: Current OCR or STS models are limited by their closed-set training that cannot handle signage recognition with arbitrary shapes, styles, and multi-view inconsistency issues in the open world (see Fig. 5). To address this, we leverage the text set of the venue maps to perform visual similarity matching that enables the closed-set detectors to adapt to the open world without fine-tuning. Specifically, we employ a two-stage process for signage recognition. In the offline stage, we utilize an off-the-shelf multilingual text-diffusion model, AnyText [34], to convert all the landmark names $\mathcal{T}$ to the text-rendered signage images set, denoted as $\{o^{\prime}\}$. Then, we adopt ESTextSpotter [28], denoted as $\phi$, to detect text regions and only extract the text features $e$ before the recognition heads from these offline-generated images without using the recognition results. This process creates a prior features pool $\mathcal{D}_{e}$. During online exploration, we utilize the same spotter $\phi$ to detect the signage from the current image observation $o_{t}$ and extract its features $e_{t}$ at timestep $t$. We then calculate the cosine similarity $s_{\phi}(o_{t},o^{\prime}_{j})$ between the detected text features and those of the generative text images to find the most similar pair. In practice, we also enhance the retrieval process by narrowing down the matching scope to a subset of $h$-hop neighboring ($NN$) to the previous recognized landmark $\tau_{t-1}$ on the venue maps based on the robot’s real-time location, resulting in better performance and faster speed:
$s_{\phi}(o_{t},o^{\prime}_{j})=cos(\phi(o_{t}),\phi(o^{\prime}_{j})),\ j\in NN(\mathcal{G},\tau_{t-1},h).$ (1)
@@
As such, the detected signage $o_{t}$ is recognized and matched to the most similar landmark name $\tau_{t}$ at the image level, i.e., $\tau_{t}=\mathcal{T}_{arg\max s_{\phi}(o_{t},o^{\prime}_{j})}$, with a recognition score $s_{\phi}(o_{t},o^{\prime}_{j})$.
@@
2D-to-3D Fusion: To mitigate the noisy text recognition from single-view 2D images, we propose adopting a 2D-to-3D instance fusion strategy [35] to enhance recognition robustness using multiple observations. During exploration, the signage map serves as a 3D text bank $\mathcal{B}$, which maintains the merged 3D text instances $R^{3D}_{j}$ and corresponding features $e_{j}^{3D}$. At each frame $o_{t}$, we first extract the 2D text features $e_{i}^{2D}$ for each detected 2D text regions $\{r_{i}^{2D},..,r_{k}^{2D}\}$, then project them to 3D regions $\{r_{i}^{3D},..,r_{k}^{3D}\}$ using depth $D_{t}$, pose $p_{t}$ and camera intrinsics $C$. Afterwards, for each new 3D text region $r_{i}^{3D}$ and each existing 3D text instance $R^{3D}_{j}$ in the bank, if their cosine similarity $cos(e^{2D}_{i},e^{3D}_{j})$ exceeds a predefined threshold $\gamma_{s}$ and the 3D distance between their centroid points $d(r_{i}^{3D},R_{j}^{3D})$ is lower than a threshold $\gamma_{d}$, then we fuse two instances by aggregating centroid points and averaging over the features:
$R^{3D}_{j}:=R^{3D}_{j}\cup r_{i}^{3D}~{}\text{and}~{}e_{j}^{3D}:=(ne_{j}^{3D}+e_{i}^{2D})/(n+1).$ (2)